OpenStudy (anonymous):

can someone help me with adding/subtracting fractions? 5 questions medal and fan

OpenStudy (comrad):

Sure thing! Just give me a little while because I'm lagging hard. Fire away with the questions though!

OpenStudy (anonymous):

so whats the question?

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

hmmm im a bit rusty but ill try

OpenStudy (comrad):

same here

OpenStudy (anonymous):

give me a sec to work it out

OpenStudy (anonymous):

I got an idea

OpenStudy (anonymous):

7/18 lemme tell you why

OpenStudy (anonymous):

it is 1/6 which becomes 3/18 + 2/9 which becomes 4/18 so 4/18 + 3/18 = 7/18

OpenStudy (comrad):

It's 7/18

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

hmm i got 3/18 idk if im right

OpenStudy (comrad):

it's 7/18

OpenStudy (anonymous):

oh ok :)

OpenStudy (comrad):

medal?

OpenStudy (anonymous):

1/6 = 3/18 and 2/9 = 4/18 so 3/18 + 4/18 = 7/18

OpenStudy (anonymous):

Comrad you can't

OpenStudy (anonymous):

you need to explain it

OpenStudy (anonymous):

it is by the rules of the site you have to show the work and math and also how to do it

OpenStudy (comrad):

well fan? lol

OpenStudy (anonymous):

I mean you have to show how you got 7/18

OpenStudy (anonymous):

umm i also pretty confused sorry, though i guess i need to study more;)

OpenStudy (anonymous):

im sorry i let you down:(

OpenStudy (anonymous):

I am serious http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

you need to explain the answer too since I got suspended once for it

OpenStudy (anonymous):

it is 1/12 since 3/4 = 9/12 and 2/3 = 8/12 so 9/12 - 8/12 = 1/12

OpenStudy (anonymous):

is this a test?

OpenStudy (anonymous):

if it isn't I can still help you

OpenStudy (anonymous):

cause openstudy

OpenStudy (anonymous):

is this a test?

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

is it a test?

OpenStudy (anonymous):

I am asking you a question that I need to know since if I don't I can get in huge trouble

OpenStudy (anonymous):

its not a test

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6 7/8

OpenStudy (anonymous):

since

OpenStudy (anonymous):

4 5/8 + 2 1/4 = 6 7/8 since 2 1/4 = 2 2/8 so 2 2/8 + 4 5/8 = 6 7/8

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

@Alicia_redd can I have a medal?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

can you post this in another tab

OpenStudy (anonymous):

by closing this tab

OpenStudy (anonymous):

Sure and sorry for the slow replies i'm lagging.